# Bifurcation sequences in the secular 3D planetary 3-Body problem: a geometric approach

### v2

@@ 4.3 Secular normal form: octupole approximation ~$\tilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$
$\Delta_{1}\!=0.000492281$ $\mathcal{D}_{2}\!=-0.330852,$ $\Delta_{2}\!=0.00187156,$
and Eqs. (40) and (41) can be solved explicitly yielding the bifurcation values in the parameter $\sigma_{0}$ where critical points of the first and second kind appear. We find $\sigma_{0}^{(CP1,1)}=0.00489265$ and $\sigma_{0}^{(CP1,2)}=0.00655611$, while the values $\sigma_{0}^{(CP2,1)}=0.00623676$ and $\sigma_{0}^{(CP2,2)}=0.00497142$ correspond to the appearance/disappearance of CP2, entering/leaving the circle $\sigma_{1}^{2}+\sigma_{3}^{2}=\sigma_{0}^{2}$ in the plane $\sigma_{2}=0$ (on the sphere, two CPs of the second kind bifurcate from/off a critical point of first kind corresponding to a tangency on the plane $\sigma_{2}=0$).
Figure 12: Phase portraits and analysis of the bifurcations (see text), same as in Figure 8, but for the Hamiltonian ~$\widetilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$.
@@
The sequence of bifurcations observed for the Hamiltonian ~$\widetilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$ can now be analyzed using the same geometric method as for the Hamiltonians $\mathcal{H}_{int}$ and $\mathcal{H}_{int}^{(1)}$ (see Fig. 12). The usual ACR states A and B are given by external tangencies between $\mathcal{S}_{\sigma_{0}}$ and $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$, the curves $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ being hyperbola-like in this case as well. The fourth panel corresponds to the bifurcation value where one of the two analytical solutions of Eq. (40) are found, $\sigma_{0}^{(CP1,2)}=0.00655611$. Note that the analytical determination of all bifurcation values is due to the simplicity of the model, which leads to analytical expressions for all coefficients. Decreasing the value of $\sigma_{0}$ (fifth panel), the phase portrait exhibits two new periodic orbits $P_{1}$ and $P_{2}$. Looking in detail the passage from the fourth to the fifth panels, the bifurcation corresponds to the fact that the green hyperbola $\mathcal{C}_{\sigma_{0}^{(CP1,2)},\,\mathcal{E}}^{(\sigma_{2}=0)}$, tangent to $\mathcal{S}_{\sigma_{0}^{(CP1,2)}}^{(\sigma_{2}=0)}$ at the point $P$ in the fourth panel, splits into two hyperbolas in the fifth panel, one (red dashed) corresponding to an outer tangency yielding the stable fixed point $P_{2}$, and the other (black dashed) corresponding to an inner tangency yielding the unstable fixed point $P_{1}$. Decreasing further the value of $\sigma_{0}$ (6th to 8th panel) the two branches of the hyperbola (black dashed) yielding the tangency at $P_{1}$ come closer and closer, until they merge at a bifurcation of a ‘CP2’ type ($\sigma_{0}^{(CP2,1)}=0.00623676$). Decreasing further $\sigma_{0}$ (9th panel) $P_{1}$ becomes stable by a pitchfork bifurcation, giving rise to two new unstable fixed points (CP2) $F_{1}$ and $F_{2}$. For still smaller $\sigma_{0}$ (10th and 11th panel) the fixed points $F_{1}$ and $F_{2}$ move away from the fixed point $P_{1}$, eventually colliding with the B mode (12th panel). This terminates the $F$-family of periodic orbits by an inverse pitchfork bifurcation which renders the point B unstable, at a value $\sigma_{0}=\sigma_{0}^{(CP2,2)}=0.00497142$ which, again, can be computed analytically through Eq. (41)). Finally, at still smaller $\sigma_{0}$ (13th panel), mode B becomes unstable (internal tangency), while, as $\sigma_{0}$ decreases, the hyperbolas passing through the tangencies at B and A (red and black in the 13th and 14th panels) approach each other and eventually collide to a single hyperbola (15th panel, cyan) at $\sigma_{0}=\sigma_{0}^{(CP1,1)}=4.89\cdot 10^{-3}$, computed also analytically. At the collision the points A and B collide to a single bifurcation point $\tilde{P}$. Then, for $\sigma_{0}<\sigma_{0}^{(CP1,1)}$ the number of tangency points passes from four to two, i.e., the only stable points surviving at the end of the bifurcation sequence are $P_{2}$ and $P_{1}$.
@@ 5 Conclusions
    Section 4 focuses on a comparison of the bifurcation sequences found in various levels of integrable Hamiltonian approximation to the full secular Hamiltonian $\mathcal{H}_{sec}$ of the planetary three body problem. A main conclusion is that considering just $\mathcal{H}_{int}$, i.e., the integrable part of $\mathcal{H}_{sec}$, leads to a sequence of bifurcations at variance to the one observed by the phase portraits of the complete secular Hamiltonian $\mathcal{H}_{sec}$. However, using a ‘book-keeping’ method in conjunction with normalization by Lie series, we arrive at a secular normal form model $\mathcal{H}^{(1)}_{int}$, which is able to qualitatively reproduce the same sequence of bifurcations as in the complete system.
@@
    A special case of the previous analysis is the one of the octupolar approximation to the Hamiltonian truncated at order 4 in the orbital eccentricities (Subsection 4.3). Computing a secular normal form as above for this last model leads to the Hamiltonian ~$\widetilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$, for which the bifurcation limits throughout the whole bifurcation sequence can be predicted analytically using the formulas of subsection 3.4. This is particularly useful in cases of so-called hierarchical systems (see, for example, [14][5] and [16]) where the octupolar expansion is sufficiently precise.
As a final remark, we emphasize that the method here discussed is generic, i.e., applicable to any integrable Hamiltonian model which admits $\sigma_{0}$ (the total angular momentum) as a second integral and approximates the Jacobi-reduced 3D secular planetary dynamics. As regards the ability to provide explicit formulas for the critical bifurcation values of the periodic orbits of the system, this is only limited by the complexity of the formulas arising as higher order truncations are considered of the initial secular Hamiltonian. Thus, in cases of hierarchical systems, where a low order truncation of the Hamiltonian model is sufficiently precise, explicit formulas are easy to provide. In non-hierarchical systems, on the other hand, predictions by the method retain essentially their analytical character, the numerical operations to be performed being limited to computing the sign-definiteness of some suitably defined quadratic forms, as well as root-finding applied to the equations defining the tangencies (points CP1) or transverse degenerate intersections (points CP2) between the surfaces $\mathcal{S}_{\sigma_{0}}$ and $\mathcal{C}_{\sigma_{0},\mathcal{E}}$.

### v3

@@ 4.3 Secular normal form: octupole approximation $\tilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$
$\Delta_{1}\!=0.000492281$ ${\mathcal{D}_{3}}\!=-0.330852,$ ${\Delta_{3}}\!=0.00187156,$
and Eqs. (47) and (50) can be solved explicitly yielding the bifurcation values in the parameter $\sigma_{0}$ where critical points of the first and second kind appear. We find $\sigma_{0}^{({CPI},1)}=0.00489265$ and $\sigma_{0}^{({CPI},2)}=0.00655611$, while the values $\sigma_{0}^{({CPII},1)}=0.00623676$ and $\sigma_{0}^{({CPII},2)}=0.00497142$ correspond to the appearance/disappearance of CPII, entering/leaving the circle $\sigma_{1}^{2}+\sigma_{3}^{2}=\sigma_{0}^{2}$ in the plane $\sigma_{2}=0$ (on the sphere, two CPs of the second kind bifurcate from/off a critical point of first kind corresponding to a tangency on the plane $\sigma_{2}=0$).
Figure 13: Phase portraits and analysis of the bifurcations (see text), same as in Figure 9, but for the Hamiltonian $\widetilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$.
@@
The sequence of bifurcations observed for the Hamiltonian $\widetilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$ can now be analyzed using the same geometric method as for the Hamiltonians $\mathcal{H}_{int}$ and $\mathcal{H}_{int}^{(1)}$ (see Fig. 13). The usual ACR states A and B are given by external tangencies between $\mathcal{S}_{\sigma_{0}}$ and $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$, the curves $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ being hyperbola-like in this case as well. The fourth panel corresponds to the bifurcation value where one of the two analytical solutions of Eq. (47) is found, $\sigma_{0}^{({CPI},2)}=0.00655611$. Note that the analytical determination of all bifurcation values is due to the simplicity of the model, which leads to analytical expressions for all coefficients. Decreasing the value of $\sigma_{0}$ (fifth panel), the phase portrait exhibits two new periodic orbits $P_{1}$ and $P_{2}$. Looking in detail the passage from the fourth to the fifth panels, the bifurcation corresponds to the fact that the green hyperbola $\mathcal{C}_{\sigma_{0}^{({CPI},2)},\,\mathcal{E}}^{(\sigma_{2}=0)}$, tangent to $\mathcal{S}_{\sigma_{0}^{({CPI},2)}}^{(\sigma_{2}=0)}$ at the point $P$ in the fourth panel, splits into two hyperbolas in the fifth panel, one (red dashed) corresponding to an outer tangency yielding the stable fixed point $P_{2}$, and the other (black dashed) corresponding to an inner tangency yielding the unstable fixed point $P_{1}$. Decreasing further the value of $\sigma_{0}$ (6th to 8th panel) the two branches of the hyperbola (black dashed) yielding the tangency at $P_{1}$ come closer and closer, until they merge at a bifurcation of a ‘ CPII’ type ($\sigma_{0}^{({CPII},1)}=0.00623676$). Decreasing further $\sigma_{0}$ (9th panel) $P_{1}$ becomes stable by a pitchfork bifurcation, giving rise to two new unstable fixed points ( CPII) $F_{1}$ and $F_{2}$. For still smaller $\sigma_{0}$ (10th and 11th panel) the fixed points $F_{1}$ and $F_{2}$ move away from the fixed point $P_{1}$, eventually colliding with the B mode (12th panel) at value $\sigma_{0}=\sigma_{0}^{({CPII},2)}=0.00497142$ which, again, can be computed analytically through Eq. (50). This terminates the $F$-family of periodic orbits by an inverse pitchfork bifurcation which renders the point B unstable. In fact, at still smaller $\sigma_{0}$ (13th panel), mode B is unstable ( having internal tangency), while, as $\sigma_{0}$ decreases, the hyperbolas passing through the tangencies at B and A (red and black in the 13th and 14th panels) approach each other and eventually collide to a single hyperbola (15th panel, cyan) at $\sigma_{0}=\sigma_{0}^{({CPI},1)}=4.89\cdot 10^{-3}$, computed also analytically. At the collision the points A and B collide to a single bifurcation point $\tilde{P}$. Then, for $\sigma_{0}<\sigma_{0}^{({CPI},1)}$ the number of tangency points passes from four to two, i.e., the only stable points surviving at the end of the bifurcation sequence are $P_{2}$ and $P_{1}$.
We conclude observing than, notwithstanding its simplicity compared to the full secular Hamiltonian $\mathcal{H}_{sec}$, the model $\tilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$ qualitatively recovers the same sequence of bifurcations as in the full secular dynamics, in the entire range of energies up to $\mathcal{E}=\mathcal{E}_{C,2}$, i.e., before the birth of the Kozai-Lidov regime. Comparing Fig. 1 and Fig. 13, despite the quantitative difference in the form of the invariant curves and the precise values of the energy where bifurcations take place, the bifurcation sequences obtained by the two models are qualitatively the same. On the other hand, quantitatively comparable results with the two models are expected to hold in the case of hierarchical systems ($a_{2}\ll a_{3}$), in which the octupolar approximation of the Hamiltonian is sufficiently accurate.
@@ 5 Conclusions
    Section 4 focuses on a comparison of the bifurcation sequences found in various levels of integrable Hamiltonian approximation to the full secular Hamiltonian $\mathcal{H}_{sec}$ of the planetary three body problem. A main conclusion is that considering just $\mathcal{H}_{int}$, i.e., the integrable part of $\mathcal{H}_{sec}$, leads to a sequence of bifurcations at variance with the one observed by the phase portraits of the complete secular Hamiltonian $\mathcal{H}_{sec}$. However, using a ‘book-keeping’ method in conjunction with normalization by Lie series, we arrive at a secular normal form model $\mathcal{H}^{(1)}_{int}$, which is able to qualitatively reproduce the same sequence of bifurcations as in the complete system.
@@
    A special case of the previous analysis is the one of the octupolar approximation to the Hamiltonian truncated at order 4 in the orbital eccentricities (Subsection 4.3). Computing a secular normal form as above for this last model leads to the Hamiltonian $\widetilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$, for which the bifurcation limits throughout the whole bifurcation sequence can be predicted analytically using the formulas of subsection 3.4. This is particularly useful in cases of so-called hierarchical systems (see, for example, [18], [6] and [20]) where the octupolar expansion is sufficiently precise.
As a final remark, we emphasize that the method here discussed is generic, i.e., applicable to any integrable Hamiltonian model which admits $\sigma_{0}$ (the total angular momentum) as a second integral and approximates the Jacobi-reduced 3D secular planetary dynamics. As regards the ability to provide explicit formulas for the critical bifurcation values of the periodic orbits of the system, this is only limited by the complexity of the formulas arising as higher order truncations are considered of the initial secular Hamiltonian. Thus, in cases of hierarchical systems, where a low order truncation of the Hamiltonian model is sufficiently precise, explicit formulas are easy to provide. In non-hierarchical systems, on the other hand, predictions by the method retain essentially their analytical character, the numerical operations to be performed being limited to computing the sign-definiteness of some suitably defined quadratic forms, as well as root-finding applied to the equations defining the tangencies (points CPI) or transverse degenerate intersections (points CPII) between the surfaces $\mathcal{S}_{\sigma_{0}}$ and $\mathcal{C}_{\sigma_{0},\mathcal{E}}$. Note that, according to figures 17 and 18 of [17], the bifurcations dealt with in the present paper appear rather commonly for generic (non-hierarchical) values of the masses and/or semi-major axes of exoplanetary systems with mutually inclined planetary orbits. For example, apart from the case of the system $\upsilon$-Andromedae studied here, a simple comparison of the masses and semi-major axes shows that all the observed systems enlisted in Table 1 of [15] have a high probability to fall in the regime of an orbital configuration corresponding to one of the possible bifurcations stemming from the basic ACR states A or B. The periodic orbits associated with such bifurcations are important also from the viewpoint that they represent the natural endstates of dissipative processes which are likely to occur in the period of formation of the exoplanetary systems.
@@
Appendix A Numerical coefficients for the integrable Hamiltonian $\mathcal{H}_{int}$
Consider the integrable Hamiltonian $\mathcal{H}_{int}$ (Eq. (5)). Up to terms of second order in the variables $\sigma_{i}$ (i.e. of fourth order in the eccentricities), apart from constants, we find the approximate formula
$\mathcal{H}_{int}\approx A\sigma_{1}^{2}+C\sigma_{3}^{2}+B\sigma_{1}\sigma_{3}+D(\sigma_{0})\sigma_{1}+E(\sigma_{0})\sigma_{3}+F(\sigma_{0})\,$,
that is equivalent to Eq. (52) of Section 4.1. For the parameters (masses, semi-major axes, and AMD value) as in the $\upsilon$-Andromedæ system (see [17]) the coefficients read:
$A\!=0.00212824,$ $C\!=0.00186469,$
$B\!=-0.00186482,$ $D(\sigma_{0})\!=0.000165361-0.0159745\,\sigma_{0},$
The coefficients $A,B,C$ satisfy $B^{2}<4AC\,$, hence the quadratic form $A\sigma_{1}^{2}+C\sigma_{3}^{2}+B\sigma_{1}\sigma_{3}$ yields ellipses.
Appendix B Numerical coefficients for the secular integrable normal form $\mathcal{H}_{int}^{(1)}$
Consider the integrable Hamiltonian $\mathcal{H}_{int}^{(1)}$ (Eq. (59)). Up to terms of second order in the variables $\sigma_{i}$ (i.e. of fourth order in the eccentricities), the secular normal form reads (apart from constants)
that is equivalent to Eq. (61) of Section 4.2. For the parameters (masses, semi-major axes, and AMD value) as in the $\upsilon$-Andromedæ system (see [17]) the coefficients read:
$A^{(1)}\!=0.00912208,$ $C^{(1)}\!=-0.0518887,$
Being $(B^{(1)})^{2}>4A^{(1)}C^{(1)}\,$, the quadratic form $A^{(1)}\sigma_{1}^{2}+C^{(1)}\sigma_{3}^{2}+B^{(1)}\sigma_{1}\sigma_{3}$ yields hyperbolas.